# Jaime Angulo Pava  and  Alexander Muñoz

### v3

@@ 4.2.1. Extensions generating contractile dynamics
Let $\textbf{U}=(\phi,\psi_{1},\dots,\psi_{N})\in D(A_{\mathcal{L}})$. From (4.7) and the definition of $A_{\mathcal{L}}$ we have
$\begin{split}2\mbox{Re}[A_{\mathcal{L}}\textbf{U},\textbf{U}]&=[A_{\mathcal{L}}\textbf{U},\textbf{U}]+\overline{[A_{\mathcal{L}}\textbf{U},\textbf{U}]}=[A_{\mathcal{L}}\textbf{U},\textbf{U}]+[\textbf{U},A_{\mathcal{L}}\textbf{U}]=-[A_{0}^{*}\textbf{U},\textbf{U}]-[\textbf{U},A_{0}^{*}\textbf{U}]\\ &=\left(\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\right)_{0}-\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\right)_{1}.\end{split}$ (4.36)
Hence, $\mbox{Re}[A_{\mathcal{L}}\textbf{U},\textbf{U}]\leqslant 0$ if and only if for all $\textbf{U}\in D(A_{\mathcal{L}})$ it follows
$\left(\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\right)_{0}\leqslant\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\right)_{1}.$ (4.37)
From (4.31) we conclude that (4.37) is equivalent to $\mathcal{L}$ being a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction. ∎
@@
Let $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$ be densely defined. $A_{\mathcal{L}}^{*}$ is dissipative if and only if $\mathcal{L}^{\#}$ is a $(\mathcal{C}_{0},\mathcal{C}_{1})$-contraction.
@@
Let $\textbf{V}\in D(A_{\mathcal{L}}^{*})$. From (4.7) we have
$\begin{split}2\mbox{Re}[A_{\mathcal{L}}^{*}\textbf{V},\textbf{V}]&=[A_{0}^{*}\textbf{V},\textbf{V}]+[\textbf{V},A_{0}^{*}\textbf{V}]\\ &=\left(\mathcal{L}^{\#}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\mathcal{L}^{\#}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)_{1}-\left(\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)_{0}.\end{split}$ (4.38)
Hence, $\mbox{Re}[A_{\mathcal{L}}^{*}\textbf{V},\textbf{V}]\leqslant 0$ if and only if
(#$\left(\mathcal{L}^{\#}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\mathcal{L}^{\#}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)_{1}\leqslant\left(\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)_{0}\ \mbox{ for all }\textbf{V}\in D(A_{\mathcal{L}}^{*}).$ (4.39)
From the definition of $D(A_{\mathcal{L}}^{*})$ we conclude (4.39) is equivalent to $\mathcal{L}^{\#}$ being a $(\mathcal{C}_{0},\mathcal{C}_{1})$-contraction. ∎
From the previous two propositions above and the Lummer-Phillips Theorem (see Theorem 2.2) we can obtain the following characterization of extensions generating $C_{0}$-semigroups of contractions.
@@
Let $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$ be densely defined. $A_{\mathcal{L}}$ is the generator of a contraction semigroup if and only if $\mathcal{L}$ is a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction and $\mathcal{L}^{\#}$ satisfies (4.39).
@@
Suppose $A_{\mathcal{L}}$ generates a semigroup $\{T(t)\}_{t\geqslant 0}$ of contractions. By Proposition 2.1, $A_{\mathcal{L}}$ is dissipative. Since $A_{\mathcal{L}}^{*}$ generates the semigroup $\{T(t)^{*}\}_{t\geqslant 0}$ which is also a contraction semigroup, from Proposition 2.1 we conclude $A_{\mathcal{L}}^{*}$ is dissipative. From Propositions 4.5 and 4.6 we conclude that $\mathcal{L}$ is a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction and $\mathcal{L}^{\#}$ satisfies (4.39). The converse follows form Propositions 4.5 and 4.6 and Theorem 2.2. ∎

### v4

@@ 4.2.2. Boundary relations, Krein adjoints, and contractive extensions
For $\mathbf{U}\in D(A_{R})$, set $(x,y):=(\Gamma_{1}\mathbf{U},\Gamma_{0}\mathbf{U})\in R$. The Green identity (4.26) applied to the pair $(\mathbf{U},\mathbf{U})$ gives
$2\operatorname{Re}[A_{R}\mathbf{U},\mathbf{U}]=(y\mid y)_{0}-(x\mid x)_{1}.$ (4.33)
Hence $\operatorname{Re}[A_{R}\mathbf{U},\mathbf{U}]\leqslant 0$ for all $\mathbf{U}\in D(A_{R})$ if and only if $(y\mid y)_{0}\leqslant(x\mid x)_{1}$ for all $(x,y)\in R$, which is exactly (4.28). ∎
Remark 2.
We note contractivity of $R$ implies contractivity of $R^{[*]}$ in this finite- dimensional setting. In fact, by the polar decomposition of an invertible Hermitian matrix we can write $C_{0}=J_{0}H_{0}$ and $C_{1}=J_{1}H_{1}$ with $H_{0},H_{1}$ positive definite and $J_{0},J_{1}$ fundamental symmetries of the respective Krein spaces. The contractivity condition (4.28) for $R$, when expressed through the similarity transforms $H_{0}^{1/2}$ and $H_{1}^{1/2}$, becomes a Hilbert-space contraction condition for an equivalent relation. Its adjoint in the Hilbert sense is again a contraction, and undoing the similarity transforms shows that $R^{[*]}$ satisfies (4.28) with the roles of $\mathcal{C}_{0}$ and $\mathcal{C}_{1}$ interchanged. Consequently the dissipativity of $A_{R}$ automatically entails the dissipativity of $A_{R}^{*}=-A_{R^{[*]}}$, without any additional hypothesis on $R$.
Theorem 4.6.
Let $R\subset\mathcal{C}_{1}\times\mathcal{C}_{0}$ be a linear boundary relation. Then $A_{R}$ generates a $C_{0}$-semigroup of contractions on $L^{2}(\mathcal{G})$ if and only if $R$ is a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction.
Proof.
Suppose $R$ is contractive. By Proposition 4.5, $A_{R}$ is dissipative. By Remark 2, $R^{[*]}$ is also contractive, so Proposition 4.5 applied to $R^{[*]}$ gives that $A_{R}^{*}=-A_{R^{[*]}}$ is dissipative. The Lumer–Phillips theorem then yields that $A_{R}$ generates a contraction semigroup.
Conversely, if $A_{R}$ generates a contraction semigroup, the Lumer–Phillips theorem implies $A_{R}$ is dissipative, and Proposition 4.5 gives that $R$ is contractive. ∎
4.2.3. Examples
We now present extensions $A_{R}$ generating semigroups of contractions, framed as boundary relations. The first example corresponds to an operator-type relation (graph of a matrix), while Remark 1 describes the additional Dirichlet-type extensions that fall outside the operator framework.
Example 4 ($\delta$-type interactions on a tadpole graph).
Let $N=1$, $\alpha_{0}=\alpha_{1}=\beta_{0}=\beta_{1}$, and let $m_{1},\dots,m_{4}\in\mathbb{R}$. Define the matrix
$M:=\begin{pmatrix}1&0&0&0&0&0\\[2.0pt] \dfrac{m_{2}+m_{3}}{2}&1&0&1&0&0\\[6.0pt] m_{1}&m_{2}+m_{3}&2&m_{2}&0&0\\[2.0pt] 0&0&0&1&0&0\\[2.0pt] \dfrac{m_{2}+m_{3}}{2}&1&0&1&0&0\\[6.0pt] m_{3}&2&0&m_{4}&0&2\end{pmatrix}$, (4.34)
and consider the boundary relation $R:=\operatorname{graph}(M)\subset\mathcal{C}_{1}\times\mathcal{C}_{0}$, i.e.
$R=\bigl\{(x,Mx):x\in\mathcal{C}_{1}\bigr\}.$ (4.35)
By Definition 4.5, $R$ is a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction if and only if the matrix $M^{*}C_{0}M-C_{1}$ is negative semidefinite. A direct computation shows that a sufficient condition is
$\alpha_{0}<0,\qquad m_{4}\leqslant\tfrac{3}{2},\qquad 4m_{1}-(m_{2}^{2}+m_{3}^{2})\leqslant 2.$ (4.36)
By Theorem 4.6, under (4.36) the operator $A_{R}$ generates a $C_{0}$-semigroup of contractions.
The Krein adjoint relation is $R^{[*]}=\operatorname{graph}(M^{\#})$ with $M^{\#}=C_{1}^{-1}M^{*}C_{0}$, given explicitly by
$M^{\#}=\begin{pmatrix}1&0&0&0&0&0\\[2.0pt] \dfrac{-m_{2}-m_{3}}{2}&\dfrac{1}{2}&0&-1&\dfrac{1}{2}&0\\[8.0pt] \dfrac{m_{1}-1}{2}&\dfrac{-(m_{2}+m_{3})}{4}&\dfrac{1}{2}&\dfrac{m_{3}}{2}&\dfrac{-(m_{2}+m_{3})}{4}&0\\[8.0pt] 0&0&0&1&0&0\\[2.0pt] 0&0&0&0&0&0\\[2.0pt] \dfrac{m_{2}}{2}&-\dfrac{1}{2}&0&\dfrac{m_{4}-1}{2}&-\dfrac{1}{2}&\dfrac{1}{2}\end{pmatrix}$, (4.37)
which is also a $(\mathcal{C}_{0},\mathcal{C}_{1})$-contraction under (4.36), in agreement with Remark 2.
Explicitly, the domain of $A_{R}$ consists of all $(\phi,\psi)\in D(A_{0}^{*})$ satisfying $(\Gamma_{1}(\phi,\psi),\,\Gamma_{0}(\phi,\psi))\in R$, that is,
$\begin{split}&\phi(-L)=\phi(L)=\psi(L),\\[4.0pt] &\phi^{\prime}(L)-\phi^{\prime}(-L)=\left(\tfrac{m_{2}+m_{3}+2}{2}\right)\phi(-L),\\[4.0pt] &\phi^{\prime\prime}(L)-2\psi^{\prime\prime}(L)=(m_{3}+m_{4})\,\psi(L)+2\,\phi^{\prime}(-L),\\[4.0pt] &\phi^{\prime\prime}(L)-2\phi^{\prime\prime}(-L)=(m_{1}+m_{2})\,\phi(-L)+(m_{2}+m_{3})\,\phi^{\prime}(-L).\end{split}$ (4.38)
In particular, setting $m_{4}=m_{1}=0$, $m_{3}=-2$, and writing $z=m_{2}/2\in\mathbb{R}$, the conditions (4.38) reduce to
$\begin{split}&\phi(-L)=\phi(L)=\psi(L),\qquad\phi^{\prime}(L)-\phi^{\prime}(-L)=z\,\phi(-L),\\[4.0pt] &\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)-\psi^{\prime\prime}(L)=z\,\phi^{\prime}(-L)+(z-1)\,\phi(-L).\end{split}$ (4.39)
The adjoint operator $A_{R}^{*}=-A_{R^{[*]}}$ has domain consisting of all $(\phi,\psi)\in D(A_{0}^{*})$ satisfying $(\Gamma_{0}(\phi,\psi),\,\Gamma_{1}(\phi,\psi))\in R^{[*]}$, i.e.
$\begin{split}&\phi(-L)=\phi(L)=\psi(L),\qquad\psi^{\prime}(L)=0,\\[4.0pt] &-\tfrac{m_{2}+m_{3}}{2}\,\phi(L)+\tfrac{1}{2}\,\phi^{\prime}(L)-\phi(-L)+\tfrac{1}{2}\,\psi^{\prime}(L)=\phi^{\prime}(-L),\\[4.0pt] &\tfrac{m_{1}-1}{2}\,\phi(L)-\tfrac{m_{2}+m_{3}}{4}\,\phi^{\prime}(L)+\tfrac{1}{2}\,\phi^{\prime\prime}(L)+\tfrac{m_{3}}{2}\,\phi(-L)-\tfrac{m_{2}+m_{3}}{4}\,\psi^{\prime}(L)=\phi^{\prime\prime}(-L),\\[4.0pt] &\tfrac{m_{2}}{2}\,\phi(L)-\tfrac{1}{2}\,\phi^{\prime}(L)+\tfrac{m_{4}-1}{2}\,\phi(-L)-\tfrac{1}{2}\,\psi^{\prime}(L)+\tfrac{1}{2}\,\phi^{\prime\prime}(L)=\psi^{\prime\prime}(L).\end{split}$ (4.40)
Remark 3.
A canonical example of extensions not reachable by linear operators is the partial Dirichlet relation
$R_{j}:=\bigl\{(x,y)\in\mathcal{C}_{1}\times\mathcal{C}_{0}:x_{j}=0\bigr\},$ (4.41)
where $x_{j}\in\mathbb{C}^{3}$ is the $j$-th block of $x$, which imposes $\partial\psi_{j}(L)=0$ (i.e. $\psi_{j}(L)=\psi_{j}^{\prime}(L)=\psi_{j}^{\prime\prime}(L)=0$) while leaving $\Gamma_{0}\mathbf{U}$ unconstrained. Such relations are contractive whenever the residual quadratic form $(y\mid y)_{0}$ is non-positive on $\{0\}\times\operatorname{mul}(R)$. No finite matrix $\mathcal{L}$ can encode these boundary conditions, since that would require $\mathcal{L}\cdot 0=y\neq 0$, which is impossible for a linear map.
4.3. Boundary system separating derivatives
A finer description of the boundary behavior is obtained by separating the quadratic form involving first derivatives from those involving the function values and second derivatives.
The technique of separating the first-derivative component from the boundary system, developed in the present subsection, is not merely a formal device. It has already proven to be an effective tool in the qualitative analysis of dispersive equations on metric graphs: in [42], this separation of derivatives was exploited to construct solitary-wave solutions for the Korteweg–de Vries equation on star graphs. Furthermore, the class of extensions $A_{Y,R_{1}}$ introduced below, together with the contraction semigroups they generate, constitutes the functional-analytic backbone of an ongoing program aimed at the stabilization and controllability of the KdV equation on metric graphs.
We recall the notation $\vec{\mathbf{U}}$, $\vec{\mathbf{U}}^{\prime}$, $\vec{\mathbf{U}}^{\prime\prime}$ introduced in Subsection 2.1.
Lemma 4.7.
For any $\mathbf{U}=(\phi,\psi_{1},\dots,\psi_{N})$ and $\mathbf{V}=(\tilde{\phi},\tilde{\psi}_{1},\dots,\tilde{\psi}_{N})$ in $D(A_{0}^{*})$,
$A_{0}^{*}\mathbf{U},\mathbf{V}]+[\mathbf{U},A_{0}^{*}\mathbf{V}]=\bigl(D_{\beta}\vec{\mathbf{U}}\mid\vec{\mathbf{V}}\bigr)+\bigl(D_{\alpha}\vec{\mathbf{U}}^{\prime\prime}\mid\vec{\mathbf{V}}\bigr)+\bigl(D_{\alpha}\vec{\mathbf{U}}\mid\vec{\mathbf{V}}^{\prime\prime}\bigr)-\bigl(D_{\alpha}\vec{\mathbf{U}}^{\prime}\mid\vec{\mathbf{V}}^{\prime}\bigr),$ (4.42)
where
$D_{\alpha}:=\operatorname{diag}(\alpha_{0},-\alpha_{0},\alpha_{1},\dots,\alpha_{N}),\qquad D_{\beta}:=\operatorname{diag}(\beta_{0},-\beta_{0},\beta_{1},\dots,\beta_{N}).$ (4.43)
Definition 4.7.
Consider the spaces $\mathcal{K}_{+}\cong\mathcal{K}_{-}\cong\mathbb{C}^{N+1}$ equipped with the weighted inner products
$(x\mid y)_{\mathcal{K}_{+}}:=(D_{\alpha}^{+}\,x\mid y),\qquad(x\mid y)_{\mathcal{K}_{-}}:=(D_{\alpha}^{-}\,x\mid y),$ (4.44)
where
$D_{\alpha}^{+}:=\operatorname{diag}(\alpha_{0},\alpha_{1},\dots,\alpha_{N}),\qquad D_{\alpha}^{-}:=\frac{1}{N+1}\operatorname{diag}(\alpha_{0},\alpha_{0},\dots,\alpha_{0}).$ (4.45)
Both $(\mathcal{K}_{\pm},(\cdot\mid\cdot)_{\mathcal{K}_{\pm}})$ are Krein spaces.
The matrices $D_{\alpha}^{-}$ and $D_{\alpha}^{+}$ split $D_{\alpha}$: the former collects the coefficient of $\phi(L)$ and the latter collects all remaining coefficients. Specifically, (4.42) and the identity
$\bigl(D_{\alpha}\vec{\mathbf{U}}^{\prime}\mid\vec{\mathbf{V}}^{\prime}\bigr)=\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\;\middle|\;\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}_{1}^{\prime}(L)\\ \vdots\\ \tilde{\psi}_{N}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}-\left(\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\;\middle|\;\begin{pmatrix}\tilde{\phi}^{\prime}(L)\\ \tilde{\phi}^{\prime}(L)\\ \vdots\\ \tilde{\phi}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}$ (4.46)
yield the split Green identity
$\begin{split}[A_{0}^{*}\mathbf{U},\mathbf{V}]+[\mathbf{U},A_{0}^{*}\mathbf{V}]&=\Bigl(D_{\alpha}\vec{\mathbf{U}}^{\prime\prime}+\tfrac{1}{2}D_{\beta}\vec{\mathbf{U}}\;\Big|\;\vec{\mathbf{V}}\Bigr)+\Bigl(\vec{\mathbf{U}}\;\Big|\;\tfrac{1}{2}D_{\beta}\vec{\mathbf{V}}+D_{\alpha}\vec{\mathbf{V}}^{\prime\prime}\Bigr)\\ &\quad-\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\;\middle|\;\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}_{1}^{\prime}(L)\\ \vdots\\ \tilde{\psi}_{N}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}+\left(\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\;\middle|\;\begin{pmatrix}\tilde{\phi}^{\prime}(L)\\ \tilde{\phi}^{\prime}(L)\\ \vdots\\ \tilde{\phi}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}.\end{split}$ (4.47)
4.3.1. Boundary relations for the first-derivative part
We introduce derivative trace maps $\Lambda_{+},\Lambda_{-}:D(A_{0}^{*})\to\mathcal{K}_{\pm}$ by
$\Lambda_{+}\mathbf{U}:=\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix},\qquad\Lambda_{-}\mathbf{U}:=\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}$. (4.48)
Definition 4.8.
A derivative boundary relation is a linear subspace $R_{1}\subset\mathcal{K}_{+}\times\mathcal{K}_{-}$. Its Krein adjoint $R_{1}^{[*]}\subset\mathcal{K}_{-}\times\mathcal{K}_{+}$ is defined by
[…]
The relation $R_{1}$ is called a $(\mathcal{K}_{+},\mathcal{K}_{-})$-contraction if
[…]
When $R_{1}=\operatorname{graph}(\mathcal{L})$ for a linear map $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$, the Krein adjoint satisfies $R_{1}^{[*]}=\operatorname{graph}(\mathcal{L}^{\#})$ with $\mathcal{L}^{\#}:=(D_{\alpha}^{+})^{-1}\mathcal{L}^{*}D_{\alpha}^{-}$, and (4.50) reduces to $(D_{\alpha}^{+})^{-1}\mathcal{L}^{*}D_{\alpha}^{-}\mathcal{L}-I\preceq 0$.
4.3.2. Extensions generating contractive dynamics via splitting derivatives
Definition 4.9.
Let $R_{1}\subset\mathcal{K}_{+}\times\mathcal{K}_{-}$ be a linear boundary relation and let $Y\subset\mathbb{C}^{N+2}$ be a closed subspace. Define
$\begin{split}D(A_{Y,R_{1}}):=\Bigl\{\mathbf{U}\in D(A_{0}^{*})\;\Big|\;&\vec{\mathbf{U}}\in Y,\quad D_{\alpha}\vec{\mathbf{U}}^{\prime\prime}+\tfrac{1}{2}D_{\beta}\vec{\mathbf{U}}\in Y^{\perp},\\ &\text{and}\quad(\Lambda_{+}\mathbf{U},\,\Lambda_{-}\mathbf{U})\in R_{1}\Bigr\},\end{split}$ (4.51)
with $A_{Y,R_{1}}\mathbf{U}:=-A_{0}^{*}\mathbf{U}$.
The subspace condition $\vec{\mathbf{U}}\in Y$ and the orthogonality condition $D_{\alpha}\vec{\mathbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\mathbf{U}}\in Y^{\perp}$ together govern how function values and second derivatives interact at the vertex. The relation condition $(\Lambda_{+}\mathbf{U},\Lambda_{-}\mathbf{U})\in R_{1}$ governs the first-derivative coupling independently.
@@ 4.3.2. Extensions generating contractive dynamics via splitting derivatives
Let $R_{1}\subset\mathcal{K}_{+}\times\mathcal{K}_{-}$ be a linear boundary relation and $Y\subset\mathbb{C}^{N+2}$ a closed subspace. Then $A_{Y,R_{1}}^{*}=-A_{Y,R_{1}^{[*]}}$, where $A_{Y,R_{1}^{[*]}}$ acts as $A_{0}^{*}$ with domain
$\begin{split}D\!\left(A_{Y,R_{1}^{[*]}}\right)=\Bigl\{\mathbf{V}\in D(A_{0}^{*})\;\Big|\;&\vec{\mathbf{V}}\in Y,\quad\tfrac{1}{2}D_{\beta}\vec{\mathbf{V}}+D_{\alpha}\vec{\mathbf{V}}^{\prime\prime}\in Y^{\perp},\\ &\text{and}\quad(\Lambda_{-}\mathbf{V},\,\Lambda_{+}\mathbf{V})\in R_{1}^{[*]}\Bigr\}.\end{split}$ (4.52)
@@
Let $\mathbf{U}\in D(A_{Y,R_{1}})$ and $\mathbf{V}\in D(A_{0}^{*})$. Using (4.47) and $A_{Y,R_{1}}\mathbf{U}=-A_{0}^{*}\mathbf{U}$,
$\begin{split}[A_{Y,R_{1}}\mathbf{U},\mathbf{V}]&=[\mathbf{U},A_{0}^{*}\mathbf{V}]-\Bigl(D_{\alpha}\vec{\mathbf{U}}^{\prime\prime}+\tfrac{1}{2}D_{\beta}\vec{\mathbf{U}}\;\Big|\;\vec{\mathbf{V}}\Bigr)-\Bigl(\vec{\mathbf{U}}\;\Big|\;\tfrac{1}{2}D_{\beta}\vec{\mathbf{V}}+D_{\alpha}\vec{\mathbf{V}}^{\prime\prime}\Bigr)\\ &\quad+(\Lambda_{+}\mathbf{U}\mid\Lambda_{+}\mathbf{V})_{\mathcal{K}_{+}}-(\Lambda_{-}\mathbf{U}\mid\Lambda_{-}\mathbf{V})_{\mathcal{K}_{-}}.\end{split}$ (4.53)
We need $A_{Y,R_{1}}\mathbf{U},\mathbf{V}]=[\mathbf{U},A_{Y,R_{1}}^{*}\mathbf{V}]$ for all $\mathbf{U}\in D(A_{Y,R_{1}})$.
Taking $\mathbf{U}$ with $\vec{\mathbf{U}}=\vec{\mathbf{U}}^{\prime}=0$ gives $(D_{\alpha}\vec{\mathbf{U}}^{\prime\prime}\mid\vec{\mathbf{V}})=0$ for all $D_{\alpha}\vec{\mathbf{U}}^{\prime\prime}\in Y^{\perp}$, hence $\vec{\mathbf{V}}\in Y$.
For arbitrary $x\in Y$, taking $\mathbf{U}$ with $\vec{\mathbf{U}}=x$ and $\vec{\mathbf{U}}^{\prime}=0$, the condition $D_{\alpha}\vec{\mathbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\mathbf{U}}\in Y^{\perp}$ and (4.53) yield
$\Bigl(x\;\Big|\;\tfrac{1}{2}D_{\beta}\vec{\mathbf{V}}+D_{\alpha}\vec{\mathbf{V}}^{\prime\prime}\Bigr)=\Bigl(D_{\alpha}\vec{\mathbf{U}}^{\prime\prime}+\tfrac{1}{2}D_{\beta}\vec{\mathbf{U}}\;\Big|\;\vec{\mathbf{V}}\Bigr)=0,$
hence $\frac{1}{2}D_{\beta}\vec{\mathbf{V}}+D_{\alpha}\vec{\mathbf{V}}^{\prime\prime}\in Y^{\perp}$.
At this point (4.53) reduces to the requirement that for all $\mathbf{U}\in D(A_{Y,R_{1}})$,
$(\Lambda_{+}\mathbf{U}\mid\Lambda_{+}\mathbf{V})_{\mathcal{K}_{+}}=(\Lambda_{-}\mathbf{U}\mid\Lambda_{-}\mathbf{V})_{\mathcal{K}_{-}}.$
Since $(\Lambda_{+}\mathbf{U},\Lambda_{-}\mathbf{U})\in R_{1}$, the definition (4.49) of the Krein adjoint shows this is equivalent to $(\Lambda_{-}\mathbf{V},\Lambda_{+}\mathbf{V})\in R_{1}^{[*]}$. ∎
We now obtain, once more, dissipativity via contractivity.
Proposition 4.9.
$A_{Y,R_{1}}$ is dissipative if and only if $R_{1}$ is a $(\mathcal{K}_{+},\mathcal{K}_{-})$-contraction.
Proof.
For $\mathbf{U}\in D(A_{Y,R_{1}})$, set $(x,y):=(\Lambda_{+}\mathbf{U},\Lambda_{-}\mathbf{U})\in R_{1}$. The split Green identity (4.47) applied to $(\mathbf{U},\mathbf{U})$, together with the conditions $\vec{\mathbf{U}}\in Y$ and $D_{\alpha}\vec{\mathbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\mathbf{U}}\in Y^{\perp}$, gives
$2\operatorname{Re}[A_{Y,R_{1}}\mathbf{U},\mathbf{U}]=(x\mid x)_{\mathcal{K}_{+}}-(y\mid y)_{\mathcal{K}_{-}}.$ (4.54)
Hence $\operatorname{Re}[A_{Y,R_{1}}\mathbf{U},\mathbf{U}]\leqslant 0$ for all $\mathbf{U}\in D(A_{Y,R_{1}})$ if and only if $(y\mid y)_{\mathcal{K}_{-}}\leqslant(x\mid x)_{\mathcal{K}_{+}}$ for all $(x,y)\in R_{1}$, i.e. $R_{1}$ is contractive. ∎
Remark 4.
The same Hilbert-space similarity argument as in Remark 2 shows that contractivity of $R_{1}$ automatically implies contractivity of $R_{1}^{[*]}$. Hence dissipativity of $A_{Y,R_{1}}$ entails dissipativity of $A_{Y,R_{1}}^{*}$.
@@
Let $R_{1}\subset\mathcal{K}_{+}\times\mathcal{K}_{-}$ be a linear boundary relation and $Y\subset\mathbb{C}^{N+2}$ a closed subspace. Then $A_{Y,R_{1}}$ generates a $C_{0}$-semigroup of contractions on $L^{2}(\mathcal{G})$ if and only if $R_{1}$ is a $(\mathcal{K}_{+},\mathcal{K}_{-})$-contraction.
@@
If $R_{1}$ is contractive, Proposition 4.9 gives dissipativity of $A_{Y,R_{1}}$, and Remark 4 gives dissipativity of $A_{Y,R_{1}}^{*}$. The Lumer–Phillips theorem yields the conclusion. Conversely, if $A_{Y,R_{1}}$ generates a contraction semigroup it is dissipative, and Proposition 4.9 gives contractivity of $R_{1}$. ∎